OpenStudy (anonymous):

simplify ( c^2 - 4cd / 3a ) ( 18a^2 b / c-4d ) how u simplify? can u show steps of canceling too pls? thank you

OpenStudy (anonymous):

multiply everything out. its fractions so you can multiply across

OpenStudy (anonymous):

No, no....cancel first.

OpenStudy (anonymous):

just a wonder estudier but how can you cancel with no common factors in the two separate fractions

OpenStudy (anonymous):

For a start, I have 18a^2b in the top and 3a in the bottom. And a (c-4d) in top and bottom.

OpenStudy (anonymous):

It's just multiplication (and the way he has typed it). U can just start cancelling right away.

OpenStudy (anonymous):

thats why i said multiply first lol...

OpenStudy (anonymous):

No need.

OpenStudy (anonymous):

\[(c^2 - 4cd)18a^2b/3a(c-4d)\]

OpenStudy (anonymous):

well no matter what you do it comes out to 6abc

OpenStudy (anonymous):

Explain it as he asked...

OpenStudy (anonymous):

U never know, u might get another medal...:-)

OpenStudy (anonymous):

i canceled and got 6abc^2 the two c's are from c^2 - 4cd

OpenStudy (anonymous):

c^2 -4 cd = c(c-4d) and you can cancel the c-4d top and bottom along with the a top and bottom and then 3/18 = 1/6.

OpenStudy (anonymous):

I have no time for people who post answers without an explanation...

OpenStudy (anonymous):

well i'll use estudier's way cause you gave me a medal :D using the equation he gave above you get 18a^2bc^2 - 72 a^2bcd ---------------------- 3ac- 12ad take out 3a your left with 6abc^2 - 24abcd --------------- c-4d take out the c-4d and your left with 6abc